# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

### v1

@@ 5 Discussion
We investigated the performance of the SDP relaxation of the K-means problem in the context of high-dimensional sparse clustering. Assuming a known isotropic covariance matrix, we derived the minimax separation bound for exact recovery of cluster labels with sparse subsets of features using SDP K-means. Our results demonstrate the critical role of variable selection in achieving exact recovery of cluster labels. Building on this insight, we proposed an iterative algorithm that integrates variable selection into the SDP relaxation framework. Our method avoids the precise estimation of nuisance parameters such as cluster centers and the covariance matrix (which may not be the identity matrix). There are two potential directions for future research: i) establish theoretical guarantees that our algorithm can identify a subset with sufficient separation, and ii) extend our minimax framework to the general case of an unknown covariance matrix.
@@ Appendix A Overview of Appendix
  • Appendix E provides additional details on the SDP objectives deferred from Sections 2.1 and 3.1, including their derivation and how they capture the sparse clustering signal.
@@ Appendix C Proof of Theorem 1
[…]
@@
Before we begin, we summarize the new notations used in the proof of Theorem 1 in Table 2 for the reader’s reference.
Table 2: Notations used in the proof of Theorem 1.
Symbol Description Type
$\Gamma_{K}$ Orthogonal complement of cluster span Subspace of $\mathbb{R}^{n}$
$\dot{\lambda}^{S}$ Constructed dual variable for trace constraint Scalar
The following statistics depend on the value of $\lambda^{S}$ and assymetric with respect to their indices:
$\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})\in\mathbb{R}^{n}$ Constructed dual variable for symmetry constraint Random vector
$\dot{\mathbf{B}}^{S}(\lambda^{S})\in\mathbb{R}^{n\times n}$ Constructed dual variable for nonnegativity Random matrix
$\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\in\mathbb{R}^{|G_{k}^{\ast}|\times|G_{l}^{\ast}|}$ $(k,l)$th off-diagonal block of $\dot{\mathbf{B}}^{S}$ Random matrix
$\dot{\mathbf{W}}^{S}(\lambda^{S})\in\mathbb{R}^{n\times n}$ A function of $\lambda^{S}$, $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$, and $\dot{\mathbf{B}}^{S}(\lambda^{S})$ Random matrix
$\mathbf{r}^{(k,l,S)}\in\mathbb{R}^{|G_{k}^{\ast}|}$ Row-sum of $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}$;   $\mathbf{r}^{(k,l,S)}\neq\mathbf{r}^{(l,k,S)}$. Random vector
$t^{S}_{lk}(\lambda^{S})$ Sum of $\mathbf{r}^{(l,k,S)}$ (not $\mathbf{r}^{(k,l,S)}$); Total sum of $\dot{\mathbf{B}}^{S}_{G_{l}^{\ast},G_{k}^{\ast}}$ Random variable
$U^{S}(\mathbf{X})$ Upper bound on $\lambda^{S}$ Random variable
$L_{1}^{S}(\mathbf{X})$ Lower bound on $\lambda^{S}$; independent of $\lambda^{S}$ Random variable
$L_{2}^{S}(\mathbf{X},\lambda^{S})$ Lower bound on $\lambda^{S}$; depends on $\lambda^{S}$ Random variable
The following statistics are random variables and asymmetric with respect to their indices:
$D^{S}_{kli}(\mathbf{X})$ Cluster-separation statistic that defines $U^{S}(\mathbf{X})$;
$I^{S}_{kl}(\mathbf{v},\mathbf{X})$ Signal–noise interaction term that defines $T^{S}_{1,kl}$ and further $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$N^{S}_{kl}(\mathbf{v},\mathbf{X})$ Noise–noise interaction term that defines $T^{S}_{2,kl}$ and further $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$T^{S}_{1,kl}(\mathbf{v})$ $I_{kl}(\mathbf{v},\mathbf{X})\cdot I_{lk}(\mathbf{v},\mathbf{X})$; Defines $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$T^{S}_{2,kl}(\mathbf{v})$ $N_{kl}(\mathbf{v},\mathbf{X})\cdot N_{lk}(\mathbf{v},\mathbf{X})$; Defines $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$T^{S}_{3,kl}(\mathbf{v})$ Mixture of $T^{S}_{1,kl}(\mathbf{v})$ and $T^{S}_{2,kl}(\mathbf{v})$; Defines $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$G^{S}_{1,kl}(\mathbf{v})$ Sum of noise inner products that comprises $N^{S}_{kl}(\mathbf{v},\mathbf{X})$
$G^{S}_{2,kl}(\mathbf{v})$ Sum of mean noise inner products that comprises $N^{S}_{kl}(\mathbf{v},\mathbf{X})$
@@ C.1.1 Necessary conditions for dual certificates
[…]
@@ C.1.3 Reformulation into high-probability bound problem
$L_{1}^{S}(\mathbf{X}):=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2},$ (20)
@@ C.2.1 High-probability Bounds
$L_{1}^{S}(\mathbf{X})=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2},$
@@ C.3.2 Bounding the signal-noise interaction term
We first provide a high-probability upper bound of
$I^{S}_{kl}(\mathbf{v},\mathbf{X})=|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;\mathcal{E}_{S\cap S_{0},i}\rangle,$
defined in (21).
@@ C.3.3 Bounding the noise-noise interaction term
$=\underbrace{\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle}_{:=G^{S}_{1,kl}(\mathbf{v})}\underbrace{-|G_{l}^{\ast}|\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,l}^{\ast},\;\mathcal{E}_{S,i}\bigr{\rangle},}_{:=G^{S}_{2,kl}(\mathbf{v})}$
@@
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}G^{S}_{1,kl}(\mathbf{v})\lesssim|G_{l}^{\ast}|\sqrt{|S|(\log n+|S|\log p)}.$
@@ C.4.1 Proof of Lemma 13
$\min_{\lambda^{S}\in\mathbb{R},\boldsymbol{\alpha}^{S}\in\mathbb{R}^{n},\mathbf{B}^{S}\in\mathbb{R}^{n\times n}}\hskip-20.00003pt\lambda^{S}K+(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{n},\quad\text{s.t.}\quad\mathbf{B}^{S}\geq 0,\mathbf{W}^{S}\succeq 0.$
@@
=$\quad\quad\stackrel{{\scriptstyle(i)}}{{=}}\lambda^{S}\mathrm{tr}(\mathbf{Z})+\frac{1}{2}(\boldsymbol{\alpha}^{S})^{\top}(\mathbf{Z}+\mathbf{Z}^{\top})\boldsymbol{1}_{n}-\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\;\mathbf{Z}\rangle$
=$\quad\quad\stackrel{{\scriptstyle(ii)}}{{=}}\bigl{\langle}\lambda^{S}\mathbf{I}_{n}+\frac{1}{2}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}+\frac{1}{2}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}-\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}-\mathbf{B}^{S},\;\mathbf{Z}\bigr{\rangle}+\langle\mathbf{B}^{S},\mathbf{Z}\rangle$
@@ C.4.2 Proof of Lemma 14
[…]
@@
=$\overset{(i)}{=}(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})_{G_{k}^{\ast},(G_{k}^{\ast})^{C}}(\mathbf{1}_{G_{l}^{\ast}})_{(G_{k}^{\ast})^{C}}$
@@
$\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}$ =$\overset{(i)}{=}\mathbf{1}_{G_{k}^{\ast}}^{\top}\biggl{\{}\lambda^{S}\mathbf{1}_{G_{k}^{\ast}}+\frac{1}{2}\bigl{(}\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\mathbf{1}_{n}+\frac{|G_{k}^{\ast}|}{2}\boldsymbol{\alpha}^{S}-(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{k}^{\ast}}-\mathbf{B}^{S}\mathbf{1}_{G_{k}^{\ast}}\biggr{\}}$
=$\overset{(ii)}{=}\lambda^{S}|G_{k}^{\ast}|+|G_{k}^{\ast}|\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}-|G_{k}^{\ast}|^{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}$ (37)
=$\overset{(iii)}{=}0,$
@@
This corresponds precisely to condition 14.
@@ C.4.4 Proof of Lemma 17
$\mathbf{v}^{\top}\dot{\mathbf{W}}^{S}\mathbf{v}$ =$\overset{(i)}{=}\lambda^{S}\mathbf{v}^{\top}\textbf{I}_{n}\mathbf{v}+\frac{1}{2}\mathbf{v}^{\top}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}\mathbf{v}+\frac{1}{2}\mathbf{v}^{\top}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}\mathbf{v}-\mathbf{v}^{\top}\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}\mathbf{v}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$
=$\overset{(ii)}{=}\lambda^{S}+(\mathbf{v}^{\top}\boldsymbol{\alpha}_{S})\sum_{i=1}^{n}v_{i}-\|\mathbf{X}_{S,\cdot}\mathbf{v}\|_{2}^{2}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$
=$\overset{(ii)}{=}\lambda^{S}-\|\mathbf{X}_{S,\cdot}\mathbf{v}\|_{2}^{2}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v},$
@@
$\lambda^{S}\geq\underbrace{\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}}_{L^{S}_{1}(\mathbf{X})}+\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}.$ (44)
@@ 4. Derivation of $T^{S}_{1,kl}$, $T^{S}_{2,kl}$, and $T^{S}_{3,kl}$
=$\quad\stackrel{{\scriptstyle(i)}}{{=}}\frac{|G_{l}^{\ast}|}{2}\sum_{i\in G_{k}^{\ast}}v_{i}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)}$
@@
=$\quad\overset{(ii)}{=}|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle\overline{\mathbf{X}}^{\ast}_{S,k}-\overline{\mathbf{X}}^{\ast}_{S,l},\;\mathbf{X}_{S,i}\rangle$
=$\quad\overset{(iii)}{=}|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S}+\overline{\mathcal{E}}_{S,k}^{\ast}-\overline{\mathcal{E}}_{S,l}^{\ast},\;{\mathcal{E}}_{S,i}\rangle$

### v2

@@ 5 DISCUSSION
Under a known isotropic covariance, we derived the minimax separation required for exact cluster recovery across varying signal feature subset estimates, achieved by a modified SDP $K$-means. This result highlights the critical role of feature selection in sparse clustering while showing that slight overselection of features is not detrimental. Building on this insight, we proposed iterative algorithms that integrate feature selection into SDP $K$-means, avoiding the need to precisely estimate nuisance parameters such as cluster centers and the covariance matrix (which may not be the identity). There are two potential directions for future research: i) establish theoretical guarantees that our algorithm can identify a subset with sufficient separation, and ii) extend our minimax framework to the general case of an unknown covariance matrix.
@@ Appendix A OVERVIEW OF APPENDIX
  • Appendix E provides additional details on the SDP objectives deferred from Sections 2 and 3.1, including their derivation and how they capture the sparse clustering signal.
@@ Appendix C PROOF OF THEOREM 1
For notational simplicity, we adopt the following matrix-based notations of our model $\mathbf{X}_{i}=\boldsymbol{\mu}_{k}^{\ast}+\boldsymbol{\varepsilon}_{i}$ for each $i\in G_{k}^{\ast}$, where $\boldsymbol{\varepsilon}_{i}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}\mathbf{I}_{p})$:
@@ C.1.1 Necessary conditions for dual certificates
$(\mathrm{C}1)\quad\mathbf{B}^{S}\geq 0;$
$(\mathrm{C}2)\quad\mathbf{W}^{S}:=\lambda^{S}\mathbf{I}_{n}-\mathbf{B}^{S}+\frac{1}{2}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}+\frac{1}{2}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}-\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}\succeq 0;$
$(\mathrm{C}3)\quad\langle\mathbf{W}^{S},\mathbf{Z}^{\ast}\rangle=0;$
$(\mathrm{C}4)\quad\langle\mathbf{B}^{S},\mathbf{Z}^{\ast}\rangle=0;$
$(\mathrm{C}5)\quad\mathbf{B}^{S}_{\mathbf{G}_{k}^{\ast},\mathbf{G}_{l}^{*}}>0\quad\text{for all}\;1\leq k\neq l\leq K.$
@@ C.1.3 Reformulation into high-probability bound problem
$L_{1}^{S}(\mathbf{X}):=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\hskip-13.00005pt\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2},$ (24)
@@ C.2.1 High-probability Bounds
$L_{1}^{S}(\mathbf{X})=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\hskip-13.00005pt\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2},$
@@ C.3.2 Bounding the signal-noise interaction term
We first provide a high-probability upper bound of $I^{S}_{kl}(\mathbf{v},\mathbf{X})=|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;\mathcal{E}_{S\cap S_{0},i}\rangle$, defined in (25).
@@ C.3.3 Bounding the noise-noise interaction term
$=\underbrace{\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\hskip-15.00002ptv_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle}_{:=G^{S}_{1,kl}(\mathbf{v})}\underbrace{-|G_{l}^{\ast}|\sum_{i\in G_{k}^{*}}v_{i}\bigl\langle\overline{\mathcal{E}}_{S,l}^{\ast},\;\mathcal{E}_{S,i}\bigr\rangle,}_{:=G^{S}_{2,kl}(\mathbf{v})}$
@@
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\hskip-10.00002ptG^{S}_{1,kl}(\mathbf{v})\lesssim|G_{l}^{\ast}|\sqrt{|S|(\log n+|S|\log p)}.$
@@ C.4.1 Proof of Lemma 13
$\min_{\lambda^{S}\in\mathbb{R},\boldsymbol{\alpha}^{S}\in\mathbb{R}^{n},\mathbf{B}^{S}\in\mathbb{R}^{n\times n}}\thinspace\lambda^{S}K+(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{n},\quad\text{s.t.}\quad\mathbf{B}^{S}\geq 0,\mathbf{W}^{S}\succeq 0.$
@@
$\quad\quad\stackrel{{\scriptstyle(i)}}{{=}}\lambda^{S}\mathrm{tr}(\mathbf{Z})+\frac{1}{2}(\boldsymbol{\alpha}^{S})^{\top}(\mathbf{Z}+\mathbf{Z}^{\top})\boldsymbol{1}_{n}-\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\;\mathbf{Z}\rangle$
$\quad\quad\stackrel{{\scriptstyle(ii)}}{{=}}\bigl\langle\lambda^{S}\mathbf{I}_{n}+\frac{1}{2}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}+\frac{1}{2}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}-\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}-\mathbf{B}^{S},\;\mathbf{Z}\bigr\rangle+\langle\mathbf{B}^{S},\mathbf{Z}\rangle$
@@ C.4.2 Proof of Lemma 14
$\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}$ $=0~\text{for }k\in[K]~(\text{by (C2) and (C3)}),$ (33)
$\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}$ $=\mathbf{0}~\text{for }k\in[K]~(\text{by (C2) and (C3)}),$ (34)
$\mathbf{B}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}$ $=\mathbf{0}~\text{for }k\in[K]~(\text{by (C4)})$ (35)
$(\mathbf{B}^{S}\mathbf{1}_{G_{k}^{\ast}})_{G_{k}^{\ast}}$ $=0~\text{for }k\in[K]~(\text{by (C4)}),\text{and}$ (36)
@@
$\overset{(i)}{=}(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})_{G_{k}^{\ast},(G_{k}^{\ast})^{C}}(\mathbf{1}_{G_{l}^{\ast}})_{(G_{k}^{\ast})^{C}}$
@@
$\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}$ $\overset{(i)}{=}\mathbf{1}_{G_{k}^{\ast}}^{\top}\biggl\{\lambda^{S}\mathbf{1}_{G_{k}^{\ast}}+\frac{1}{2}\bigl(\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr)\mathbf{1}_{n}+\frac{|G_{k}^{\ast}|}{2}\boldsymbol{\alpha}^{S}-(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{k}^{\ast}}-\mathbf{B}^{S}\mathbf{1}_{G_{k}^{\ast}}\biggr\}$
$\overset{(ii)}{=}\lambda^{S}|G_{k}^{\ast}|+|G_{k}^{\ast}|\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}-|G_{k}^{\ast}|^{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}$ (41)
$\overset{(iii)}{=}0,$
@@
This corresponds precisely to condition 18.
@@ C.4.4 Proof of Lemma 17
$\mathbf{v}^{\top}\dot{\mathbf{W}}^{S}\mathbf{v}$ $\overset{(i)}{=}\lambda^{S}\mathbf{v}^{\top}\textbf{I}_{n}\mathbf{v}+\frac{1}{2}\mathbf{v}^{\top}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}\mathbf{v}+\frac{1}{2}\mathbf{v}^{\top}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}\mathbf{v}-\mathbf{v}^{\top}\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}\mathbf{v}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$
$\overset{(ii)}{=}\lambda^{S}+(\mathbf{v}^{\top}\boldsymbol{\alpha}_{S})\sum_{i=1}^{n}v_{i}-\|\mathbf{X}_{S,\cdot}\mathbf{v}\|_{2}^{2}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$
$\overset{(ii)}{=}\lambda^{S}-\|\mathbf{X}_{S,\cdot}\mathbf{v}\|_{2}^{2}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v},$
@@
$\lambda^{S}\geq\underbrace{\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\hskip-13.00005pt\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}}_{L^{S}_{1}(\mathbf{X})}+\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\hskip-13.00005pt\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}.$ (48)
@@
$\quad\stackrel{{\scriptstyle(i)}}{{=}}\frac{|G_{l}^{\ast}|}{2}\sum_{i\in G_{k}^{\ast}}v_{i}\bigl(\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr)$
@@
$\quad\overset{(ii)}{=}|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle\overline{\mathbf{X}}^{\ast}_{S,k}-\overline{\mathbf{X}}^{\ast}_{S,l},\;\mathbf{X}_{S,i}\rangle$
$\quad\overset{(iii)}{=}|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S}+\overline{\mathcal{E}}_{S,k}^{\ast}-\overline{\mathcal{E}}_{S,l}^{\ast},\;{\mathcal{E}}_{S,i}\rangle$